Spring 2015 6GeV Commissioning Plan

From Hall D Ops Wiki
Jump to: navigation, search

This page gives the revised run plan for the Spring 2015 commissioning based on a 5.5GeV beam. It is a controlled document that can be edited only by E. Chudakov, E. Smith, A. Deur or David Lawrence. Suggestions and comments should be send to them. This supersedes the original Spring 2015 Commissioning Plan.

Items in red indicate missing information.

Items in pink indicate information requiring verification.

General Information

This document describes the run plan and procedures for the Hall D spring 2015 commissioning with 5.5 GeV beam.

  • Schedule for the Spring 2015 run. (Dates are according to the Schedule. They will be updated if necessary in the run coordination meeting pages)
    1. April 8- April 14: Electron/Photon beam restoration (6 days)
    2. Apr. 15 - May 4: Physics run (Halls A/B/D).
  • Goals for the Spring 2015 run.
  1. Commission the new Hall D equipment:
    • LH2 target, high priority. ( Expected time: ?? (Estimated 28h for 10.5 GeV Beam) )
    • Diamond radiator/goniometer, high priority. ( Expected time 80h )
    • Accelerator RF timing, low priority. ( Expected time ?? )
    • Total Absorption Counter (TAC), low priority. ( Expected time ?? )
    • Photon beam polarimeter, low priority. ( Expected time ?? )
  2. Establish that we can run with Solenoid at 1300A (implies ramping it up to 1350A).
  3. Establish DAQ operation in Block Mode with ~ 20 kHz rate, high priority. ( Expected time 16h )
  4. Complete detector calibration/alignment, high priority. ( Expected time ?? )
  5. Complete trigger optimization, high priority. ( Expected time 80h )
  6. Commission beam stability feedbacks. ( Expected time a day ?? )
  7. Test diamond on 3.4mm hole, low priority. ( Expected time ?? )
  8. πo calibration data for both FCAL and BCAL ( Expected time 144h )
  9. Chamber alignment data ( Expected time 20h )
  • Expected Staffing and responsibilities:
    • The Run Coordinator (RC, Alexandre Deur) oversees the commissioning.
    • The Physics Division Liaison (PDL, Benedikt Zihlmann) and Hall D Work Coordinator/Safety Warden (T. Carstens) verify that the proper safety rules are followed.
    • The Analysis Coordinator (TBD) will organize the off-line analysis effort.
    • Individuals responsible for each commissioning step are named in the section describing the step. They are also responsible for logging in the ELog the summary of how the step proceeded and other relevant information.
    • Shifts
      • 2 persons per shift, one with solenoid training
  • Coordination Activities:
    • The RC will attend the MCC meetings at 7:45am and 8:00am and the weekly MCC meeting (Wednesday, 1:30pm). The PDL will attend the weekly MCC meeting.
    • Daily run meetings broadcasted on Bluejeans will occur at 8:45am in the counting house conference room. The Run Coordinator will be responsible for organizing and chairing the meeting.
  • Useful links:

Beam Configuration Parameters

  • Beam energy: expected: 5.5 GeV
  • Solenoid current up to 1300A (running condition).
  • Beam current for Photon beam operations: CW 50 na-1 μA. 250 MHz frequency.
  • Radiators:
    • Nominal: Diamond, 3 test diamonds (J1a50, J2a100 S145-S90) and 3 test foils (two 10 μm Al foils, and one 100 μm Al foil with a holed 10 μm Al foil pressed on it (3mm hole)).
    • Other available:
      • 1.5 µm Al (1.7·10-5 R.L.)
      • 10 µm Al (11.2·10-5 R.L.)
      • 30 µm Al (33.7·10-5 R.L.)
  • Collimator hole: 5 mm diameter nominally. May run on the 3.4 mm once priority plans are completed.
  • Main target: LH2 cryogenic target
    • Backup targets are are:
      • 10mm thick disk (3.19 cm diameter) of CH2 (HDPE, high density polyethylene, density 0.93-0.97 g/cm3 ), 0.95 g/cm2, ~1.2% R.L.
      • 2mm thick disk (3.19 cm diameter) CH2 (HDPE)
      • 2mm thick inside a 20mm diameter area and 7mm thick outside (HDPE)
      • 3.4mm thick graphite (3.19 cm diameter) disk, 0.75 g/cm2, 1.8% R.L.
      • Al. cross made of 1.59 mm diameter wires
      • 1cm long Al. cylinder with an equilateral triangle hole inside. The triangle edge lengths are 2.5 cm

Run Plan


We expect a very limited amount of beam time (approx. 2 weeks) and will run a reduced program. There are some activities like trigger and goniometer optimization that will require multiple shifts, but also experts that will need to spread those shifts over multiple days. Therefore, the basic plan will be to establish a data taking configuration for πo running that can be used overnight to collect data and then use the days to do trigger and goniometer work.

Restore Electron Beam

  • Accelerator responsibility.
  • If Hall D tagger is used as a dump for CW beam during electron beam restoration then
    • Hall D RC must be notified.
    • Radiator, both amorphous AND diamond, must be retracted. Collimator should be in blocking position.
    • Radiation levels should be closely monito#ff00ff by MCC crew.
      • Typical levels are (for radiator retracted, collimator fully blocking, electron beam current 50 nA)
        • RAD102_P1 ~ 3 mrad/hr (tagger area, gammas, between tagger and dump)
        • RAD102_P2 ~ 15 mrad/hr (tagger area, gammas, near electronics racks) This probe is now shielded and hence should see less radiations than during the Fall 2014 run.)
        • RAD102_P3 ~ 0.005 mrem/hr (tagger area, neutrons, near electronics racks)
        • RAD101_P1 < 0.01 mrad/hr (collimator cave, gammas). This probe is now shielded and hence should see less radiations than during the Fall 2014 run.
    • Hall D tagger CARMS, Ion Chamber and BLM threshold for beam trips should not be changed without approval of Hall D leader.

During beam restoration, Accelerator will commission our beam position/energy locks (slow feedback at about 1Hz). (Note: Locks requires a minimal current 30nA).

Restore Photon Beam. Expected time: 16h

  1. Prerequisites: Amorphous AND diamond radiators retracted. Collimator should be in blocking position. Electron beam is restored and radiation levels are acceptable. Pair Spectrometer detectors/field on and its DAQ running Target status: Empty
  2. Call MCC to inform them we will Insert 2*10^-5 radiator. Insert the radiator. Check the beam position on the profiler. Adjust it if necessary. x and y should be 0±0.5cm. If such adjustments are necessary, remind MCC that they should use only the last corrector (5C11A) to adjust the beam position. If the adjustment results in unacceptable e- beam position in the tagger dump, remind MCC that they should use the tagger magnet only to readjust the e- beam position in the tagger dump.
  3. Insert the 5mm collimator. Check beam position on the Profiler and Active collimator. Check that radiation levels are good. Run for 7min to gather enough data for radiation level studies.
    • Typical levels are (for 2*10^-5 radiator, electron beam current 50 nA)
      • RAD102_P1= 3-10 mrad/hr (tagger area, gammas, between tagger and dump)
      • RAD102_P2= ?10-20? mrad/hr (tagger area, gammas, near electronics racks) This probe now has a different shielding than for Fall 2014)
      • RAD102_P3= 0.500 mrem/hr (tagger area, neutrons, near electronics racks)
      • RAD101_P1= 12.0 mrad/hr (collimator cave, gammas).
  4. Hall A leakage study.
    • Make sure Hall A is running, ask MCC for which laser they are using and the opening of their slit.
    • Ask MCC to close our slit. Insert 3*10^-4. Record rates in radiation monitors and profiler (wait there 5 min).
    • If we see any rates above background, repeat previous step with 10^-4 radiator.
    • If we still see any rates above background, repeat previous step with 2*10^-5 radiator.
    • Repeat previous step with 2*10^-5 radiator.
  5. Radiation level studies (0.5h):
    • Run 7 min without radiator (Call MCC and ask to mask the FSD each time the radiator is moved). Beam current: 50nA
    • Run 7 min with the 10^-4 radiator (Call MCC). Beam current: 50nA
    • Run 7 min with the 3*10^-4 radiator (Call MCC). Beam current: 50nA
    • Insert the 2*10^-5 radiator (Call MCC).
  6. Determine maximum photon beam transmission by doing 2D scans (x: beam scan, y: beam scan. ±5cm with 0.5cm steps). Meanwhile, monitor the rate at the active target and rates (including coinc.) in the pair spectrometer.

Verify subsystems under beam condition. Expected time: 8 h

Target is nominally empty. Solenoid at 1200A (may be less if magnet experts deem so). If the rate is not large enough with the thin radiator, put the 3.4 10^-5 radiator in and increase current to 100 nA beam. Photon should be beam optimized to max. transmission.

  1. Trigger
  2. DAQ
  3. GlueX/Tagger detectors
    • TAGH & TAGM
    • ST
    • FCAL
    • BCAL (Include Threshold scans)
    • TOF
    • FDC
    • CDC
  4. Online monitoring
    • (Pair Spectrometer): should be already running

Microscope bias Voltage study Expected time: 1.5h

Run 30 min at 3 different bias voltage values. (A. Barnes)

Commission cryogenic target. Expected time: ?? (Estimated 28h for 10.5 GeV Beam)

  • Alignment verification. Expected time: ?? (Estimated 26h for 10.5 GeV Beam)
    • Pre-requisite: Target is cold and empty, 3.4 10^-5 radiator in, 100 nA beam. Photon beam optimized to max. transmission. Solenoid at 1200A. Gluex detectors operational.
    1. If needed, empty the target: 15min
    2. Scan the beam horizontally with collimator:scan between ±20mm around nominal 111.7 position. Do 2mm steps. Spend 1min on each steps. Make sure the P.S. DAQ is running and the general DAQ is running. Correlate rates from S.T. and current in FDC with collimator position to verify target horizontal alignment. P.S. rate must normalize the ST and FDC rates due to the scan through the non-flat beam profile (if the fast feedback from the active target is enabled, this correction should be minimized). 2h.
    3. Run  ?? shifts to gather data to verify (Simon T.) that:
      • The distances between the various windows on the beam line are correct.
      • The ratio of the rate from each window to the solid target at the nose of the ST are correct.

*Density check (optional, 2h).

    1. Fill the target (30 min).
    2. Run with the target not subcooled, 30 min at 3 radiator thicknesses, 50 nA.
    3. Subcool the target (30min).
    4. Run with the target subcooled, 30 min at 3 radiator thicknesses.
  • Verify that rates scale with radiator thickness × e- beam current.

Trigger optimization. Expected time: 10 shifts spanning 12 days. Part of the commissioning should happen during the ramp up/down of the solenoid e.g. for FDC alignment

  • PS trigger with tagger (1 shift)
  1. Tune trigger parameters (timing offsets), prepare DAQ trigger configuration (no steady beam required)
(take calibration data with PS/tagger crates, 6 hours of steady beam)
  • FCAL and BCAL triggers (3 shifts)
  1. Perform threshold scans, tune timing offsets (steady beam required)
  2. Tune parameters for the FCAL & BCAL trigger
  • Commission ST/TOF triggers (1 shift)
  • Commission coincidence triggers for BCAL/FCAL and ST/TOF (2 shift)
- Trigger for FDC alignment
- Study FCAL coincidence with ST
  • Commission multiple triggers (3 shifts)
  1. Tune trigger delays and prescaling factors (steady beam required)
  2. Integrate PS, prepare trigger configuration files

DAQ tests. (2 shifts spanning 2 days)

  1. Integrate PS into DAQ readout (2h)
    • Physics trigger (already exists)
  2. Take data in various configurations (10 min. each) recording event rate, data rate, and livetime
    • mode 8: all detectors
    • mode 7: all detectors
    • mode 8: FDC, mode 7: all others
    • mode 8: CDC, mode 7: all others

High solenoid current run (1 day, concurrently to other studies)

  1. Run a day with Solenoid at 1300A

Test goniometer control. Expected time: 12 shifts spanning 12 days

Part of the plan can be done during solenoid ramp up/down.

  • --- Part 1 : Initial setup and alignment (90 μm diamond) --- (3 shifts of 8 h, not back to back). Hovanes/Ken L.
    • Prerequistes:
      • Warn in advance MCC that during this procedure, they will be able to determine the Ion Chamber trip limits for the various foils/diamonds of the goniometer. (So their appropriate personnel should be around.)
      • Stable beam is available (Locks and FFB working)
      • Tagger and microscope scalers working
      • Working PS scalers desirable
      • Main triggers / PS triggers can be run during these tests, but are not needed for the coherent brem. setup.
    • Call MCC to tell them we will scan the beam position on the goniometer to identify the foils/diamonds positions. Beam will go through the goniometer frame so the Ion Chambers should be masked.
    • Test beam position with empty holder and foils (Hovanes).
    • Tell MCC to unmask Ion Chambers.
    • Tell MCC that they can now determine the Ion Chamber trip limits for each of the foils/diamonds. This will require us to move each of them into beam position (30min).
    • Scan beam on diamond (Hovanes/Paul M.). First, use JD70-3. Then J1a50 and S145-S90. Check the difference between the three diamonds.
  • Once the diamond is aligned, take data with the PS trigger (10M triggers with the diamond radiator and PS trigger). These data must be taken during a period with stable beam.

  • --- Part 2 : commission goniometer--- (approx. 9 shifts spread over 11 days)
    • Prerequisites:
      1. Stable beam (lock commissioned and on).
      2. 5 nA of beam current. MCC needs to be warned several days in advance to be able to deliver 5nA (Note: Locks requires a minimal current 30nA. Fast Feedback requires more (unknown value for now).
      3. tagger hodoscope, microscope, photon and PairSpec scalers beam profiler operational.
    • Ken Livingston will be at JLab between Apr. 17th to Apr 30th.
      • systematic studies of scanning and debugging of the tools for running and setting up coherent brem. during Ken L. 4 night shifts 19-22 April.
    • When available, install production diamonds (50 μm), preferably during opportunistic access: (8h)

πo calibration data: Expected time: Expected time: 6 days (minimum, depends on radiator x current)

  1. BCal: Run a shift at lower temperature (Elton) 8h.
    • Controlled access necessary to change the temperature. Assume it will be opportunistic.
    • Regular trigger. Trigger mode 7.
  2. \pi ^{0} calibration data. (based on Elton's calorimeter report, here.)
    • Take 6 days of good data with FCal. Assuming 50% general efficiency, this should take about 2 weeks.
    • Concurrently, take at least 3 days of good data with BCal.
    • Take 1h run in mode-8 every day. Other runs are in mode-7, so that about 90% of the data are in mode-7.
    • Trigger type ?
  3. Run on diamond radiator with 3.4mm collimator hole (1 days?)
    • Request controlled access. Switch Active collimator from 5mm hole position to 3.4mm hole position (2h). Alex Barnes
    • Re-establish photon beam beam on 3.4mm collimator hole. Test Active collimator signals (2h). Alex Barnes
    • Take tagger and PS data in this configuration. Alex Somov.

Forward Drift Chamber Alignment. Expected time: 20h

  • Prerequisite: FCal+TOF trigger tested and optimized.
  • Conditions: B=0, 2×10-5 amorphous radiator. I<100nA Beam current, trigger type:FCal+TOF
  1. Bring down the magnetic field to 0. Ideally, it should be done opportunistically, e..g during beam studies, and determine when we will start this task: (6h)
  2. Take data (Lubomir): 20h
  3. Bring the magnetic field back up. 6h, opportunistically.

Commission fast feedbacks. Expected time 8h The four 2h tests can be done when the solenoid is ramping up/down

  • Fast Feedback (1kHz feedback) will not be ready before mid-April
  • Fast Feedback requires a minimal current of probably more than 50nA.
  • Twice, 2 hours to test the magnets and BPM/AC response. (Trent Allisson)
  • Twice, 2 hours to calibrate and turning on FFB. (Trent Allisson)

Commission Accelerator RF timing signal.

Data will be taken during concurrently with other parts of the commissioning.</strike>

==Commission Total Absorption Counter. 8h== Prerequisites: Warn MCC several days in advance that we will want to run at 5nA. (Note: Locks requires a minimal current 30nA. Fast Feedback requires more (unknown value for now).

  • Stable low current beam established (I<5nA)
  • Specific DAQ setup. Trigger: TagM, TagH. DAQ must include TAC and PS.


    1. Establish a stable low current beam <5nA: 4h
    2. Take two 2h runs: 4h

Commissioning time chart

Commissioning time chart, assuming 10 days of beam. Time t=0 corresponds to the Hall being locked-up for beam tuning. The horizontal scale is the time during which beam is available (in other word, it would be the real time with 100% of beam efficiency).